# TermSight: Making Service Contracts Approachable

### v1

@@ Abstract.
Terms of Service (ToS) are ubiquitous, legally binding contracts that govern consumers’ digital interactions. However, ToS are not designed to be read: they are filled with pages of ambiguous and complex legal terminology that burden potential users. We introduce TermSight, an intelligent reading interface designed to make ToS more approachable. TermSight offers visual summaries that highlight the relevance and power balance of information in a ToS. TermSight also categorizes and simplifies information within the ToS into concise plain-language summaries. To aid in reading the original text, TermSight offers contextualized definitions and scenarios for unfamiliar phrases. Our within-subjects evaluation of TermSight (N=20) revealed that TermSight significantly reduced the difficulty of reading ToS and increased participants’ willingness to do so. We also observed emerging strategies that participants took when interacting with AI-powered features that highlight the diverse ways that TermSight assisted ToS reading.
large language models, legal documents, reading support tools
copyright: acmlicensedjournalyear: 2025ccs: Human-centered computing Interactive systems and tools
Figure 1. The TermSight system provides guidance and reading support at multiple levels of ToS reading. At the contract level, TermSight visualizes the relevance and power balance of content in each policy (1). At the document level, TermSight chunks, summarizes, and categorizes content into 1-sentence plain-language summaries (Summary Snippets) highlighted with colors that reflect power and relevance (2). Readers can click the Summary Snippet (2) or Highlight Bar (3) to navigate between the summary snippets on the left and the original text on the right. At the phrase level, TermSight offers phrase definitions and scenarios for unfamiliar and vague phrases (4).
@@ 1. Introduction
Terms of Service (ToS)—the ubiquitous, rarely-read but legally binding contracts that govern the use of digital services—are widely incomprehensible to potential consumers. Despite consumers’ self-reported concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007), and their eventual regret when they encounter issues related to this information (e.g., the selling of their data)  (Good et al., 2005; Obar and Oeldorf-Hirsch, 2020), almost no one reads ToS content fully  (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020). There is good reason not to do so: ToS are not designed to be read. ToS are filled with legal jargon that can derail general-audience (also referred to as ‘lay-audience’) readers with confusing or under-defined terms (e.g., a ToS stating the service will collect certain ‘information’)  (Ibdah et al., 2021; Tang et al., 2021; Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). ToS also contain more than one policy (e.g., a privacy policy) that in turn each contain reams of boilerplate or semi-boilerplate text. Within these pages of text, there might be information relevant to a reader (e.g., an arbitration clause or copyright policy) but such information is lost among the many pages of text  (Srinath et al., 2021; Obar and Oeldorf-Hirsch, 2020). Legal scholars have argued that these ambiguous, jargoned, and exceptionally long contracts are designed not to cultivate a shared understanding, but to mitigate liability and establish unilateral control (Kar and Radin, 2019; Benoliel and Becher, 2019; Karanicolas, 2021; Palka, 2023).
Summarizing and simplifying text is a first step in making ToS more approachable, but resolves only part of the challenge these contracts pose. Prior interfaces have leveraged crowdsourcing (TOS, 2024; pri, 2024) and, more recently, AI-based approaches like large language models (LLMs) to extract and process text automatically from ToS (Pan et al., 2023). However, these approaches primarily focus on presenting information from pre-defined categories (e.g., user privacy) or on single policies within a ToS (e.g., a privacy policy). In reality, ToS contain many policies that cover different information (§3.3.1), and what information a reader needs is dependent on their planned usage of the site (e.g., a seller vs. buyer on an e-commerce site).
We envision a new reading experience for ToS. Our goal is to enable readers to find information important to them anywhere within a ToS and scaffold their ability to understand and contextualize this information within their potential use of a service. We start by developing a more nuanced understanding of readers’ information needs and challenges when reading ToS through a formative observational study (N=20). We found that participants cared about different information based on their envisioned usage of a service (e.g., a content creator will have different needs than a passive consumer on a social media website) and their own personal values (e.g., people had different levels of comfort when it came to privacy and data sharing). When participants sought to resolve these information needs, they encountered barriers at every level of granularity within a ToS. Within the full ToS contract, participants struggled to navigate nested policies and decide which policies to read. Within a specific policy of the ToS (e.g., the privacy policy), participants struggled to surface and interpret relevant information from extended, visually dense, and difficult text. When reading the text itself, participants struggled to interpret jargon and phrases that were ambiguous or vague. Interestingly, existing affordances (e.g., policy names, section headers with table of contents, and summaries) were often not helpful and at times misleading. For example, participants often found that policy names and section headers were unclear and that overview summaries used deceptive friendly language to hide unfavorable clauses.
We address the specific challenges surfaced by our formative work with a new intelligent reading interface, TermSight. TermSight provides multi-level reading guidance and support through contract-level overviews (Power Meter), document-level plain-language summaries (Summary Snippets), and phrase-level definitions with scenarios of potential implications (Phrase Scope). To evaluate TermSight, we conducted a counterbalanced within-subject study (N=20) comparing TermSight and a baseline HTML reader during a timed reading task. Participants found reading ToS to be significantly easier with TermSight and were more willing to read and spend time on ToS. Participants reported that features of TermSight provided guidance and reading support at all granularities of ToS, resolving challenges surfaced in our formative study. We also observed interesting and unexpected user behaviors centered around trade-offs between TermSight’s AI-generated features.
In this paper, we make the following contributions:
@@
    We characterize the challenges readers face when trying to make sense of ToS contracts. Readers face challenges at all granularities of a ToS, from determining which policies to read to being able to resolve individual ambiguous terms within the text.
@@
    We design and develop an intelligent reading interface, TermSight, for helping readers make sense of ToS contracts at the contract, document, and phrase levels.
@@
    We collect empirical evidence from our user study (N=20) demonstrating the potential value of the interaction affordances of TermSight.
@@ 2. Background and Related Work
2.1. Service Contracts and The Duty to Read
A Terms of Service (ToS) is a type of standard form contract consisting of a body of policies and conditions outlined by a service provider, dictating the rules and expectations for both the service provider and the consumer. These contracts have evolved to be more legally comprehensive compared to earlier ToS  (Amos et al., 2021), encompassing a wide range of provisions such as copyright, privacy, returns, acceptable use, and service-specific terms. By clicking agree to a ToS, a legally binding contract is formed  (Dickens, 2007). U.S. contract law is premised on the doctrine of the ‘duty to read’ where contracting parties are legally responsible for reading and understanding the contract  (Benoliel and Becher, 2019). However, the ‘duty to read’ was established in the era when contracts were short and based on mutual negotiation  (Kar and Radin, 2019). Modern ToS often have important terms buried within a large body of text and documents  (Srinath et al., 2021). The language used to write ToS can also require advanced reading capability  (Srinath et al., 2021; Amos et al., 2021) and contain unfamiliar (Ibdah et al., 2021; Tang et al., 2021) or ambiguous terminology (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). Due to the cognitive load associated with reading ToS, most consumers do not read modern ToS (Obar and Oeldorf-Hirsch, 2020; Bakos et al., 2014), despite the binding nature of ToS and consumers’ self-reports indicating concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007). ToS can also vary in content (Liu et al., 2021a; Fiesler et al., 2016) and structure (Ibdah et al., 2021) across services, leading to false consumer expectations. This widespread disengagement highlights what Kar and Radin described as a ”paradigm slip” in contract law, where agreements once premised on shared understanding and mutual consent are now reduced to unilateral boilerplate ”pseudo-contracts”, eroding fundamental principles of contract law and personal autonomy  (Kar and Radin, 2019; Palka, 2023; Benoliel and Becher, 2019). Our formative study complements prior findings by aggregating and uncovering consumers’ varying information needs and additional challenges encountered when navigating and reading policies within a ToS.
@@ 2.2. Augmenting Terms of Service
There are two major threads of work on improving the readability of ToS that we cover below: changing how ToS are designed or presented (§2.2.1) and providing summaries of ToS (§2.2.2).
2.2.1. Stylistic Manipulations
@@
To help readers better understand the implications of using a service, prior work has investigated design recommendations for service providers to make documents within ToS more accessible  (Kay and Terry, 2010; Habib et al., 2021; Taber et al., 2020). Kay and Terry (2010) proposed Textured Agreements that use typographic manipulation, pull quotes, vignettes, and iconic symbols to improve reader attention and comprehension on privacy policy compared to plain text. Habib et al. (2021) investigated how icons and linked text could be designed to better convey privacy choices. Taber et al. (2020) proposed crowdsourced sentiment highlighting of sentences in ToS, providing visual cues to guide readers toward sentences in ToS with negative sentiment. Design recommendations on styling user agreements have also been proposed  (Kay and Terry, 2010; Habib et al., 2021). While helpful in navigating contracts, these design changes do not simplify the content itself and do not take into consideration the fact that readers have different information needs when reading a ToS.
2.2.2. Summarization
@@
Work has also proposed helping consumers gain an overview of specific policies within a ToS, often the privacy policy (PP), without requiring reading the original document. Platform for Privacy Preferences (P3P) is an early attempt to standardize privacy policies by allowing service providers to submit privacy policies in a machine-readable format  (Reagle and Cranor, 1999). Tools such as Privacy Bird build atop the P3P format to provide warnings when a site’s PP does not match consumers’ privacy preferences  (Cranor et al., 2006). However, P3P is rarely adopted by service providers  (Cranor et al., 2008) and no longer supported  (Microsoft, 2016). Work has also proposed the use of a single-page summary for end-user license agreements (EULAs)  (Good et al., 2005, 2007), comic-based summaries for privacy policies  (Tabassum et al., 2018), and Privacy Policy Nutrition Labels (PPNL) that label privacy policies in a table format  (Kelley et al., 2009; Reinhardt et al., 2021) to help users gain an overview of a policy.
@@
Many earlier tools required buy-in from service providers to design better ToS. Other work, though, has focused on providing an overview summary of a ToS automatically or via crowdsourcing. For example, ToS;DR leverages volunteers to label and summarize terms in the privacy policy and the home page of ToS  (TOS, 2024). Works like CLAUDETTE automatically detect a predefined set of potentially unfair clauses from ToS  (Lippi et al., 2019; Guarino et al., 2021; Braun and Matthes, 2021; Braun et al., 2017). PrivacyCheck (Zaeem et al., 2018; Nokhbeh Zaeem et al., 2020, 2022), PrivacyGuide (Tesfay et al., 2018), and Polisis  (Harkous et al., 2018) extract terms relevant to a list of privacy-related questions and criteria and generate a grading for each. Sathyendra et al. extract opt-out choices from privacy policies  (Sathyendra et al., 2017; Bannihatti Kumar et al., 2020). Relatedly, some works evaluate the completeness of privacy policies  (Costante et al., 2012) and the GDPR compliance of extracted clauses (Liu et al., 2021b). More recently,  Pan et al. (2023) investigated generating privacy nutrition labels from a privacy policy. Work has also attempted to generate an aggregated summary based on the extracted information from a privacy policy  (Tomuro et al., 2016; Keymanesh et al., 2020). TermSight extends this rich prior work by combining the benefit of plain language summaries with interactive affordances that enable readers to navigate the original ToS text, helping readers navigate and read the original body of policies within ToS beyond privacy policies.
2.3. Intelligent Reading Interfaces
TermSight takes inspiration from a rich body of work on intelligent reading interfaces that support readers in navigating documents and clarifying information. Prior work has explored helping readers navigate documents and locate relevant information in scholarly documents. For example, work has highlighted key information in research papers, such as objectives, results, and methods, to support skimming  (Fok et al., 2022). Also in research papers, ScholarPhi identifies position-sensitive definitions of terms and symbols, enabling readers to track their meaning across contexts  (Head et al., 2021). Talk to Papers allows readers to ask questions about research papers directly, while Qlarify supports clarification by letting readers expand text spans in paper abstracts to generate explanations  (Zhao and Lee, 2020; Fok et al., 2024a). In medical research papers, Paper Plain provides a list of guiding questions to help readers navigate important information in a medical paper (August et al., 2023). In the business domain, Marco allows readers to search and ask questions across collections of business documents  (Fok et al., 2024b). In the news domain, Chen et al. (2023) introduced Marvista, a reading interface that identifies the most summative portions of news articles based on a reader’s time constraints or topic in the news article.
Compared to prior reading interfaces that have focused on research papers, news articles, or business documents, TermSight focuses on helping general audience readers make sense of legal documents, namely, ToS. Legal documents, and specifically ToS, present challenges different from prior reading contexts. In contrast to other knowledge-intensive settings (e.g., scientific or medical text), general audience readers in the legal setting are faced with language not designed to communicate information. Readers then struggle to (1) determine what text to read within and across policies, (2) define and contextualize vague language, and (3) navigate existing affordances that are themselves ambiguous or incomplete. Prior intelligent reading interfaces and context often assume that there is a single document (August et al., 2023; Fok et al., 2022; Chen et al., 2023), readers have consistent information needs  (August et al., 2023), readers know what they need to read or ask  (Fok et al., 2022; Zhao and Lee, 2020; Fok et al., 2024b), or readers can make sense of term definitions  (Head et al., 2021; August et al., 2023; Fok et al., 2024a). These existing affordances, while powerful, do not fully support readers’ needs and challenges encountered when reading digital contracts such as ToS. TermSight takes inspiration from prior works to explore the design of reading interfaces to help readers surface and interpret relevant information from legal documents.

### v2

@@ Abstract.
Legal contracts govern much of our society, but their specialized language is difficult for non-experts to read. While AI has enabled simplification of complex language, legal contracts pose unique challenges because of their connection to readers’ values, ambiguity, and legally binding nature. Based on a formative study (N=20) using Terms of Service (ToS) as example contracts to study challenges in contract reading, we developed TermSight, an intelligent reading interface to probe the opportunities and challenges of designing augmentations for legal text. TermSight guides readers to relevant clauses with color-coded plain-language snippets of information and contextualizes ambiguous language with definitions and hypothetical scenarios. Importantly, TermSight’s features always foreground the original, legally-binding contract text (e.g., linking to associated clauses). Our within-subjects study (N=20) demonstrated the opportunities of TermSight in making ToS significantly easier to read and navigate while revealing the challenges of augmenting service contracts such as ToS.
Legal Contracts, Augmented Reading, Large Language Models
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791387isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Applied computing Law
Figure 1. TermSight provides multi-level support for reading Terms of Service (ToS). At the contract level, TermSight visualizes the relevance and power balance of content in each policy (1). At the document level, TermSight chunks, summarizes, and categorizes content into one-sentence plain-language summaries (Summary Snippets) highlighted with colors that reflect power and relevance (2). Readers can click the Summary Snippet (2) or Highlight Bar (3) to navigate between the two. At the phrase level, TermSight offers phrase definitions and hypothetical scenarios for unfamiliar and ambiguous phrases (4).
The TermSight system provides guidance and reading support at multiple levels of Terms of Service (ToS) reading. The interface contains three panels: a top navigation panel for switching between different policies, the original text on the right, and summaries on the left. (1) At the contract level, TermSight visualizes the distribution of content in each policy in the top navigation panel as a horizontal bar plot based on its relevance to the user and power balance (favoring tenant vs. landlord). (2) Within each policy, TermSight chunks, summarizes, and categorizes content into one-sentence plain-language summaries (Summary Snippets), highlighted with colors that reflect power (red = favor provider, yellow = neutral, green = favor user) and relevance (low saturation = low relevance, high saturation = high relevance). The summaries are shown in the left panel, with corresponding highlight bars aligned to the original text on the right. (3) Readers can click either a Summary Snippet or its Highlight Bar to navigate between the summary and the source text. (4) At the phrase level, TermSight provides definitions and scenarios for unfamiliar or ambiguous terms.
@@ 1. Introduction
Legal contracts govern much of our society. We sign contracts for where we live, who we work with, and, increasingly, for our digital interactions online. In order to encompass many possible situations and actions, contracts contain highly specialized language, such as:
“You grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content.” — Reddit Terms of Service  (85)
On the one hand, this specialized language allows contracts to articulate the entitlements, obligations, and prohibitions (Sancheti et al., 2023) of the signing parties. On the other hand, this specialized language is exceptionally difficult for many people without a legal background to read  (Furth-Matzkin, 2017; Mueller, 1970; Amos et al., 2021).
Recent AI systems, driven by language models (LMs), have expanded our ability to transform and augment text documents for different readers. Past reading systems have simplified scientific papers (Shin et al., 2024), medical research (August et al., 2023), and news articles (Chen et al., 2023) for general audience readers. These transformations have also begun to be applied to legal language, by, for example, generating summaries based on predefined categories of rights and responsibilities in contracts (Sancheti et al., 2023; Pan et al., 2023) or through explaining legal concepts (Jiang et al., 2024).
However, legal text, and specifically contracts, pose unique challenges for simplification that current systems fail to address. First, contracts are inherently value-dependent: they are intended to encode and protect the values of the signing parties (Haapio and Passera, 2017; Fiesler et al., 2016; Sancheti et al., 2023), which may lead to varying information needs among individuals. Second, legal language is not just complex, but can be ambiguous (e.g., “We may share some of your personal information…”  (Bhatia et al., 2016)). While this language provides the linguistic flexibility to manage unforeseen circumstances  (Zhang, 2015), it also derails general audience readers  (Ibdah et al., 2021; Tang et al., 2021; Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). Finally, the literal language of a contract is important: it is the only language that is legally binding  (Cheong et al., 2024). Consequently, any text transformations should support rather than replace reading the original text.
In this paper, we explore augmentations that account for the unique challenges of simplifying legal contracts. We focus on Terms of Service (ToS), the ubiquitous but rarely-read contracts that govern the use of digital services (Obar and Oeldorf-Hirsch, 2020; Good et al., 2005; Ahmad et al., 2020). Despite consumers’ concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007), these contracts remain difficult to engage with. While prior work often focuses on a single policy within the ToS (e.g., a privacy policy  (Nokhbeh Zaeem et al., 2022)), we explore challenges readers face when trying to engage with the full ToS contract. In a formative study (N=20) where participants read ToS, we found that these contracts represented the extremes of the difficulties that legal contracts pose more generally (i.e., value-dependent, ambiguous, and legally binding), validating this initial focus. The parts of the ToS that were important to participants depended heavily on their envisioned usage and personal values, and this value-dependent information was lost among the many nested policies. ToS also contained complex and ambiguous terms that confused readers. Finally, existing features to support contract reading (e.g., policy names and section headers) often did not accurately represent the actual text and were not legally binding. [^1] ToS often include clauses such as “Headings are used in these Terms for reference only and will not be considered when interpreting them”  (85).
In response, we designed a novel intelligent reading interface for ToS contracts, TermSight. In contrast to prior work that provides overview summaries detached from the original text  (99; 83; S. Pan, T. Hoang, D. Zhang, Z. Xing, X. Xu, Q. Lu, and M. Staples (2023)), we explore augmentations that support contract reading while acknowledging the unique characteristics of legal text. TermSight provides visual overviews of policies within a ToS (Power Meter, Figure 1.1), guides readers to relevant clauses with color-coded plain-language snippets of information (Summary Snippets, Figure 1.2), and contextualizes unfamiliar or ambiguous language with definitions and scenarios of potential implications (Phrase Scope, Figure 1.4). Importantly, TermSight’s features always foreground and link to the original legal text.
To explore the opportunities and challenges of designing contract reading interfaces, we conducted a counterbalanced within-subjects study (N=20) using TermSight and a baseline HTML reader. Compared to prior focus on simplified reading tasks (e.g., reading a single clause (Robinson and Zhu, 2020; Lee et al., 2025) or policy overview (Tabassum et al., 2018; Korunovska et al., 2020)), participants engaged with real ToS contracts with 10+ policies. In the study, participants found reading ToS to be significantly easier with TermSight while also being more willing to read the original ToS. Participants reported that TermSight provided multi-level guidance and support while letting them drill down into clauses within the ToS important to them. We also observed that participants often used TermSight’s generated text (e.g., Summary Snippets) as an entrance to, or a verification of, the original ToS text. At the same time, our findings revealed the broader limitations of technological solutions like TermSight on facilitating a deeper understanding of ToS due to the overwhelming amount of potentially relevant information. We end by discussing the implications of our findings for inspiring future contract interfaces and regulatory innovation. In summary, this paper makes the following contributions:
@@
    We characterize the challenges of contract reading, specifically ToS, including determining which policies to read, resolving ambiguous terms, and navigating misleading organization within the text.
@@
    We develop a novel intelligent reading interface, TermSight, to explore the design opportunities and challenges for helping readers engage with ToS while focusing on the original, legally binding language.
@@
    We collect empirical evidence from our user study demonstrating the values of TermSight’s features in finding relevant information, navigating ambiguous or complex language, and associating generated text with original text, while revealing the challenges of improving understanding of ToS.
@@ 2. Background and Related Work
2.1. Service Contracts
A Terms of Service (ToS) is a type of standard form contract consisting of a body of policies and conditions outlined by a service provider, dictating the rules and expectations for both the service provider and the consumer. While prior work in HCI mainly focused on single policies within the ToS (e.g., a privacy policy  (Nokhbeh Zaeem et al., 2022)), these contracts encompass a wide range of provisions such as copyright, privacy, returns, acceptable use, and service-specific terms  (Amos et al., 2021). Important terms are often buried within this large body of documents  (Srinath et al., 2021). ToS can vary in content (Liu et al., 2021; Fiesler et al., 2016) and structure (Ibdah et al., 2021) across services. The language used to write ToS can also require advanced reading capability  (Srinath et al., 2021; Amos et al., 2021) and contain unfamiliar (Ibdah et al., 2021; Tang et al., 2021) or ambiguous terminology (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). Most consumers do not read modern ToS (Obar and Oeldorf-Hirsch, 2020; Bakos et al., 2014), despite the binding nature of ToS (Dickens, 2007) and consumers’ self-reports indicating concern for the information presented in ToS  (Milne and Culnan, 2004; Norberg et al., 2007). This widespread disengagement highlights what Kar and Radin described as a “paradigm slip” in contract law, where agreements once premised on shared understanding and mutual consent are now reduced to unilateral boilerplate “pseudo-contracts”, eroding fundamental principles of contract law and personal autonomy  (Kar and Radin, 2019; Palka, 2023; Benoliel and Becher, 2019). Our formative study further supplements prior findings by revealing the challenges readers face when navigating the entire ToS and the presence of misleading navigational affordances (§3).
@@ 2.2. Augmenting Terms of Service
Prior work has investigated new designs for readers navigating ToS  (Kay and Terry, 2010; Habib et al., 2021; Taber et al., 2020). Kay and Terry (2010) proposed Textured Agreements that use typographic manipulation, pull quotes, vignettes, and iconic symbols to improve reader attention and comprehension of privacy policy compared to plain text. Habib et al. (2021) investigated how icons and linked text could be designed to better convey privacy choices. Taber et al. (2020) proposed crowdsourced sentiment highlighting of sentences in ToS. Design recommendations on styling user agreements have also been proposed  (Kay and Terry, 2010; Habib et al., 2021).
@@
Work has also explored helping consumers gain an overview of specific policies within a ToS, often the privacy policy (PP), without requiring reading the original document. Platform for Privacy Preferences (P3P) was an early attempt to standardize privacy policies by allowing service providers to submit privacy policies in a machine-readable format  (Reagle and Cranor, 1999). Tools such as Privacy Bird build atop the P3P format to provide warnings when a site’s PP does not match consumers’ privacy preferences  (Cranor et al., 2006). However, P3P was not widely adopted by service providers  (Cranor et al., 2008) and is no longer supported  (Microsoft, 12/15/2016). Work has also proposed the use of a single-page summary for end-user license agreements (EULAs)  (Good et al., 2005, 2007), comic-based summaries for privacy policies  (Tabassum et al., 2018), and Privacy Policy Nutrition Labels (PPNL) that label privacy policies in a table format  (Kelley et al., 2009; Reinhardt et al., 2021).
@@
Other work has focused on providing overview summaries automatically or via crowdsourcing. ToS;DR uses volunteers to label and summarize terms in the privacy policy and the home page of ToS  (99). Works like CLAUDETTE automatically detects a predefined set of potentially unfair clauses from ToS  (Lippi et al., 2019; Guarino et al., 2021; Braun and Matthes, 2021; Braun et al., 2017). PrivacyCheck (Zaeem et al., 2018; Nokhbeh Zaeem et al., 2020, 2022), PrivacyGuide (Tesfay et al., 2018), and Polisis  (Harkous et al., 2018) extract terms relevant to a list of privacy-related questions and criteria and generate a grading for each. Past work has relied on either service providers or volunteers to make stylistic changes to contracts or automated tools to provide a summary external to the contract. In this paper, we explore new ways of combining these two threads to enable readers to navigate and read the original ToS made possible by the powerful text transformation capabilities of language models.
2.3. Augmented Reading Interfaces
Prior work has explored augmenting other types of documents, such as research papers (Lo et al., 2024) and news articles (Chen et al., 2023), with generative features. Work has provided faceted highlighting in research papers to support skimming  (Fok et al., 2023) and context-sensitive definitions of terms and symbols  (Head et al., 2021) to support reading. Also in research papers, Qlarify (Fok et al., 2024a) enabled readers to expand paper abstracts with generative explanations. Systems have also provided automated question answering for research papers (Zhao and Lee, 2020) and medical papers (August et al., 2023). Shin et al. (2024) transformed research papers into summaries of design implications. Newman et al. (2024) developed tools that synthesized multiple research papers to help researchers conduct literature reviews. In the business domain, Marco allows readers to search and ask questions across collections of business documents  (Fok et al., 2024b). In the news domain, Chen et al. (2023) introduced Marvista, a reading interface that identifies the most summative portions of news articles.
In contrast to other documents explored previously, contracts—and specifically ToS—pose unique challenges to readers. Contracts directly connect with a reader’s personal values: one reader might prefer stronger privacy safeguards while another might care about licenses  (Haapio and Passera, 2017; Fiesler et al., 2016; Sancheti et al., 2023). Contract language is also meant to be interpreted; this means there are often ambiguous terms or incomplete information (e.g., the definition of ‘information’ when discussing personal information collection), with the assumption that the reader has the requisite knowledge to fill in missing information  (Bhatia et al., 2016). Finally, contract language itself is the only legally binding text (Cheong et al., 2024), meaning any text transformations must maintain and display a tight connection with the original text. We take inspiration from past augmented reading systems to investigate what features can support readers in the unique context of contract reading.